OpenStudy (anonymous):

to the nearest tenth of a cubic centimeter what is the volume of the sphere r=7cm

OpenStudy (anonymous):

(4*pi*r^3)/3 =1436.03

OpenStudy (kenny):

Use the formula, V = (4/3) (pi) r^3. Hence, V=(4/3)(pi)(7^3)= 1436.75504 rounded to 1436.8 cubic centemeters if using a calculator w/ a pi button. That's the trick in the problem: If you use abbreviated version of 3.14 for pi, you'll end w/ 1436. 026 which rounds to 1436 to the nearest tenth. Take care to use the value for pi suggested by the teacher.